$SPAD/src/input functioncode.input

Timothy Daly

December 3, 2016

Abstract

The command

)set system functioncode on

will emit the generated lisp code when a function is compiled.

Contents

2

— * —

)set break resume
)sys rm -f functioncode.output
)spool functioncode.output
)set message auto off
)clear all

--S 1 of 3
)set system functioncode on
--R
--E 1
--S 2 of 3
a(x) == x + 1
--R
--R                                                                   Type: Void
--E 2
--S 3 of 3
a(3)
--R
--R   Compiling function a with type PositiveInteger -> PositiveInteger
--R Generated LISP code for function:
--I(|*1;a;1;frame0|
--R    (LAMBDA (|#1| |envArg|)
--I      (SPADCALL |#1| 1 (ELT |*1;a;1;frame0;MV| 0))))
--R
--R   (2)  4
--R                                                        Type: PositiveInteger
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4